OpenStudy (anonymous):

I need help remembering simple algebra. Can someone explain how to evaluate for the problem below.

OpenStudy (penguin1234):

what type of algebra problem

OpenStudy (anonymous):

Evaluate the following expression for m=36 \[\sqrt{4m} +5\]

OpenStudy (anonymous):

im thinking we do \[\sqrt{4*36}\] = \[\sqrt{144} = 12 \] and then 12 -5 = 7 but I don't think im doing this right

OpenStudy (loser66):

Yes, it is not right, It is strongly correct. hihihihihi

OpenStudy (anonymous):

wait, what? so I did it right?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

No, your adding 5 not subtracting...

OpenStudy (loser66):

oh ya!! mistypo there, should be +

OpenStudy (anonymous):

thank you! I didn't realize I had messed up on the sign giving me the wrong ending answer, but thank you!

OpenStudy (anonymous):

No problem.